OpenStudy (anonymous):

Help! I give medals:) Pic below find the height and surface area of the right circular cylinder with the height equal to the diameter. Show work

OpenStudy (anonymous):

Height=2r Volume-128pic

OpenStudy (anonymous):

Surface area of a circle is \[r ^{2} \pi \] Volume is surface area of a circle multiplied by height. So, 128 = 2r * r (squared)*P

OpenStudy (anonymous):

how do I solve for r?

OpenStudy (anonymous):

\[r ^{2}\pi * 2r = 182\] \[r ^{3}\pi=64 \] \[r = \sqrt[3]{20.37}\] r=2.73

OpenStudy (anonymous):

OOh ok LOL Thank You so much :)

OpenStudy (anonymous):

Numbers are weird, but that's the only way I see how it's done :)

OpenStudy (anonymous):

as long as the answer n work r right that's all that matter in math LOL